let P be the point (1,0) and Q a point on the curve y2=8x. the locus of the midpoint of PQ is

Grade:12th pass

1 Answers

Deepak Kumar Shringi
5 years ago
Q (2t2,4t) and P (1,0)
so mid point of PQ is h = (1+2t2)/2........(i) and k = (4t+1)/2 so t= (2k-1)/4 put in
equation 1 to get locus